# Conformal Prediction for Time-series Forecasting with Change Points

### v3

@@ Abstract
Conformal prediction has been explored as a general and efficient way to provide uncertainty quantification for time series. However, current methods struggle to handle time series data with change points — sudden shifts in the underlying data-generating process. In this paper, we propose a novel Conformal Prediction for Time-series with Change points (CPTC) algorithm, addressing this gap by integrating a model to predict the underlying state with online conformal prediction to model uncertainties in non-stationary time series. We prove CPTC’s validity and improved adaptivity in the time series setting under minimum assumptions, and demonstrate CPTC’s practical effectiveness on 6 synthetic and real-world datasets, showing improved validity and adaptivity compared to state-of-the-art baselines. Our code is open source at https://github.com/huiwenn/CPTC.
@@ 1 Introduction
Uncertainty Quantification (UQ) is a key building block of reliable machine learning systems. Conformal prediction (CP) is a popular distribution-free UQ method that provides finite sample coverage guarantees without placing assumptions on the underlying data generating process (vovk2005algorithmic, ; lei2018distribution, ). Because of its simplicity and generality, conformal prediction has seen wide adoption in many tasks, ranging from healthcare angelopoulos2022conformal , robotics lindemann2023safe , to validating the factuality of generative models cherian2024large .
@@
State Space Models (SSMs) (kalman1960new, ; aoki2013state, ) are a powerful tool for modeling time series data, as they provide a structured framework to capture the underlying temporal dependencies through latent states. In particular, Switching Dynamical Systems (SDS) (ackerson1970state, ) is a type of SSM with an additional set of latent variables (known as switches) that represent the operating mode active at the current timestep. SDS introduces the inductive biases that allow SSMs to switch between a discrete set of dynamics, which can be learned from training data. It is shown to be a flexible, robust, and interpretable representation of time series with varying dynamics exhibiting change points (ansari2021deep, ; chen2022learning, ). SDS is deployed not only for energy karakatsani2008intra ; song2014short ; fan2008machine and traffic yu2004switching ; cetin2006short forecasting as examples given above, but also on a wide range of applications including neuroscience glaser2020recurrent ; rabinovich2006dynamical ; kato2015global , engineering strachan2011measuring ; wang2007actively , and sports analytics yamamoto2018switching , all would benefit from robust and calibrated uncertainty quantification.
@@ 2 Related Work
Probabilistic forecasting has become central to modern time-series analysis by modeling a distribution over future outcomes rather than a single-point forecast (benidis2022deep, ). Classic approaches often adopt Bayesian or approximate Bayesian strategies for uncertainty quantification: for instance, Bayesian Neural Networks (BNNs) (wang2019deep, ; wu2021quantifying, ; gal2017concrete, ) or neural process models (garnelo2018conditional, ; sun2024data, ) use neural networks to parameterize the underlying stochastic process. Other methods, such as DeepAR (salinas2020deepar, ), PatchTST (nie2023patchtst, ), or Temporal Fusion Transformer (lim2021temporal, ), generate probabilistic forecasts directly by minimizing a pinball loss (i.e., quantile loss), thereby learning predictive distributions.
A key challenge arises when change points occur - i.e., abrupt distribution shifts in the time series (montanez2015inertial, ). Change point detection and segmentation have been studied extensively (beyond our scope, see (aminikhanghahi2017survey, ) for a recent survey). Modeling time series with change points typically features state-space models (SSM), and in particular switching dynamical systems (SDS) (johnson2016composing, ; ansari2021deep, ; liu2023graph, ). These SDS approaches use variational inference to fit neural parameterizations of transitions and emission distributions. However, existing probabilistic forecasters often lack strict calibration guarantees and can be overconfident in practice (lakshminarayanan2017simple, ), and many require specialized architecture tuning for different tasks, limiting their applicability to real-world UQ challenges.
@@
Conformal prediction (CP) (vovk2005algorithmic, ) has emerged as a leading framework for UQ due to its algorithmic simplicity, broad applicability, and finite-sample coverage guarantees; see angelopoulos2021gentle for a comprehensive introduction. For time-series forecasting, some works focus on joint coverage over fixed-length horizons (Schaar2021conformaltime, ; sun2023copula, ; zhou2024heterogeneous, ); others assume stable underlying dynamics and adapt only to shifts in residual distributions (xu2021conformal, ) . tibshirani2019conformal uses known covariate shifts (by propensity weighting) to achieve theoretical coverage, but can fail in settings where the shift mechanism is unknown, which is typical for time series forecasting. hajihashemi2024multi introduces a multi-model ensemble CP framework, optimizing model selection within a pool of experts for combined coverage. CP has also been extended for change-point detection in time series via conformal martingales (vovk2021retrain, ; volkhonskiy2017inductive, ).
Recent works on online conformal prediction develop online strategies that guarantee marginal coverage for adversarial or nonstationary data streams. ACI (gibbs2021adaptive, ), AgACI (zaffran2022adaptive, ), DtACI (gibbs2024conformal, ), and Conformal PID control (angelopoulos2024pid, ) leverage online optimization and have a ”reaction” period when distribution shifts. Multivalid Prediction (MVP) (bastani2022practical, ) provides group coverage guarantees and learns a calibration threshold online, although it can require longer calibration windows. Closer to our work, SPCI (xu2023sequential, ) learns an autocorrelative estimation of the residual’s quantiles through Quantile Random Forest; whereas HopCPT (auer2023hopcpt, ) condition their conformal interval on similar parts of the time series using a Modern Hopfield Network. The drawback of these two methods is that the regression models are trained during inference time, which (1) requires a long cold start window to learn meaningful associations and (2) is computationally expensive. Our contribution is a CP algorithm that is computationally light during inference time, and can leverage state prediction capabilities within the SDS model to anticipate and adapt quickly to distribution changes.
@@ 3.1 Conformal Prediction
We say a sample is covered if the true value lies in the prediction interval $y^{n+1}\in\Gamma^{1-\alpha}(x^{n+1})$ Conformal prediction is guaranteed to produce valid prediction intervals (vovk2005algorithmic, ) if the calibration data and test data are exchangeable (in a dataset $\{(x^{i},y^{i})\}_{i=1}^{n}$ of size $n$, any of its nn! permutations are equally probable).
@@ 3.2 Switching Dynamical Systems for Modeling Time-series with Change Points
There have been many different variations and implementations of SDS. Classic SDS have memory-less state transitions, i.e. $p(z_{t}|z_{t-1})$ is parameterized by a stochastic transition matrix $\mathbf{A}\in\mathbb{R}^{K\times K}$. This results in an “open loop” system and the state duration follows a geometric distribution. Recent works have sought to fix this issue, for example by closing the loop using recurrent dependencies on $s_{t}$ (linderman2016recurrent, ), or modeling explicit durations (ansari2021deep, ). Our generative model is illustrated in figure 2): for generality, we factor out any non-Markovian covariates to model parameters $\theta$. The joint distribution can be factorized as
@@ 17 $\Gamma_{t}(x_{t})\leftarrow$ Aggregate the $\Gamma_{z,t}(x_{t})$s by Eqn 10 or Eqn 11;
The exchangeability assumption does not apply to the online setting of our work. If we naively calibrate with online data, when the data distribution changes, the coverage probability may deviate from the target level $1-\alpha$. Therefore, we follow online CP methods such as ACI (gibbs2021adaptive, ) to address this problem by continuously updating $\alpha_{z,t}$ to track a surrogate miscoverage rate $\alpha_{z,t}^{*}$ as defined in eqn 8 (for each state $z$ in our setting), an internal estimate of the target coverage.
@@
In eqn 8 $M_{z,t}$ measures the probability that the true label $y_{t}$ falls outside the prediction set $\Gamma_{z,t}$ when the predicted state is $z$. We track $\alpha^{*}_{z,t}$ over time by updating the estimate $\alpha_{z,t}$ with online optimization. In our implementation, we use the simple update from ACI (Eqn 9), though more sophisticated approaches such as (angelopoulos2024pid, ) may be used in its place without affecting the overall algorithm.
@@
Our algorithm can be viewed as running multiple instances of online conformal inference (lines 14-16 of algorithm 1), one for each underlying state, and adaptively aggregating the confidence intervals when the underlying mode is switching. This differs from the setting where the temporal correlation between residuals needs to be learned online (gibbs2024conformal, ; bastani2022practical, ; xu2023sequential, ; auer2023hopcpt, ); the state model allows us to leverage training data for such information. Our algorithm is modular - the state model, forecaster model, and the online adaptive conformal prediction algorithms all operate independently of each other. For example, the state-specific online coverage tracking (line 15 of algorithm 1) can be replaced with other online conformal prediction variants for data-specific applications. Our theoretical guarantees are agnostic to the forecaster models used.
@@ 5 Experiments
Table 1: Performance on synthetic and real-world datasets with target confidence $1-\alpha=0.9$ (for horizon $T=200$ for synthetic datasets, $300$ for electricity and traffic, and $60$ for bee, mean $\pm$ standard deviation of the test samples). Methods that are invalid (coverage below $90\%$) are grayed out. Our method achieves a high level of calibration (coverage is close to $90\%$) consistently.
@@
We selected the following baseline methods. RED-SDS (Recurrent Explicit Duration Switching Dynamical System) ansari2021deep is a state-of-the-art Bayesian model that learns state transitions (implementation details in Appendix B.1). We also use REDSDS as the base predictor for mode switching for our method in the experiments. CP is a generalization of conformal prediction to the online setting, also known as Online Sequential Split Conformal Prediction in zaffran2022adaptive . We choose ACI (gibbs2021adaptive, ) to represent online CP algorithms leveraging online optimization. SPCI (xu2023sequential, ) and HopCPT auer2023hopcpt are state-of-the-art CP algorithms that learns adaptive predictive intervals by quantile regression and Modern Hopfield Networks learned from the time series respectively.
@@
For real-world datasets, the Electricity and Traffic datasets from (Dua2017UCI, ) have hourly frequency and exhibit seasonality both in terms of the time series itself and volatility. The honey bee trajectory dataset (oh2008learning, ) is the most complex, composed of 4-dimensional trajectories with length averaging to 900 frames, where the bees’ dance can be decomposed into “left turn”, “right turn” and “waggle”.
@@ 5.1 Analysis of Coverage and Sharpness Results
Results for uncertainty quantification in table 4 show that we achieve significantly better validity on all datasets compared to baselines. RED-SDS’s intervals fail due to lack of calibration, a common pitfall of probabilistic methods (shown also in xu2023sequential ). CP’s under-coverage shows that the data are not exchangeable. In the pretense of change points, ACI’s coverage fluctuates dramatically and did not converging within the horizon (more visualizations in Appendix B.5). For SPCI and HopCPT, which also guarantees asymptotic validity, the invalidity is likely because they are designed for large datasets ($T\geq 10,000$ for HopCPT), and did not have enough data to react or learn useful residual patterns for the relatively short prediction horizon of our datasets ($T=200$ for synthetic datasets, 300 for electricity and traffic, and 60 for bees). Figure 3 provides a qualitative example of how the prediction intervals compare across different methods. Our method (purple shaded regions) does not over-cover during nonvolatile hours, nor under-cover during busy ones, as does ACI.
@@ A.2 Proofs: Lemma
We derive the following lemma from the convergence result of Proposition 4.1 in [23] with two additional assumptions also according to [23]: (1) the existence of a single fixed optimal target $a^{*}\in[0,1]$, and (2) $\mathbb{E}[err_{t}|a_{t}]=M(a_{t})$. These are reasonable assumptions that allow us to quantify distribution shifts and are necessary for analyzing adaptation behavior (whereas proposition 4.1 itself is a worst-case upper bound.) We refer the readers to the original paper for the complete derivation.
@@ B.3 Dataset descriptions, continued
Electricity and Traffic. The Electricity and Traffic dataset are from the UCI machine learning repository [19] and we use the train/test split from GluonTS [2]. The electricity dataset consists of hourly electricity consumption data for 370 customers, each with 5,833 observations. The traffic dataset consists of hourly occupancy rates from 963 road sensors, each with 4,000 data points.
Dancing Bees. The honey bee trajectory dataset is from [40]. Honey bees convey information about the location and distance to a food source through a dance carried out inside the hive. This dance consists of three distinct phases: “left turn”, “right turn”, and “waggle”. The duration and orientation during the waggle phase represent the distance and direction to the food source. Experiments are conducted on 6 trajectories, with lengths of 1058, 1125, 1054, 757, 609, and 814 frames, respectively. In this paper, the prediction interval is on the first two dimensions (the coordinates of the bee) only, but all dimensions are used for prediction.
@@ B.4 Additional Experiments
We add comparison to three more baselines that handles distribution shift in time series data. They are DtACI [5], ECI [55], and L-ARC [61].
@@
Table 4: Performance on synthetic and real-world datasets with target confidence $1-\alpha=0.9$ (for horizon $T=200$ for synthetic datasets, $300$ for electricity and traffic, and $60$ for bee, mean $\pm$ standard deviation of the test samples). Methods that are invalid (coverage below $90\%$) are grayed out. Our method achieves a high level of calibration (coverage is close to $90\%$) consistently.
@@
We designed the number of latent states to be a hyperparameter given by the user. Our method ensures that the validity guarantee still holds even if the state cardinality is incorrect (theorem 4.2), as incorrect state cardinality can be formulated as a case of incorrect classification. Determining state space cardinality is a well-studied problem (see for example [14]), and we consider it to be beyond the scope of our paper.

### v4

@@ Abstract
Conformal prediction has been explored as a general and efficient way to provide uncertainty quantification for time series. However, current methods struggle to handle time series data with change points — sudden shifts in the underlying data-generating process. In this paper, we propose a novel Conformal Prediction for Time-series with Change points (CPTC) algorithm, addressing this gap by integrating a model to predict the underlying state with online conformal prediction to model uncertainties in non-stationary time series. We prove CPTC’s validity and improved adaptivity in the time series setting under minimum assumptions, and demonstrate CPTC’s practical effectiveness on 6 synthetic and real-world datasets, showing improved validity and adaptivity compared to state-of-the-art baselines. Our code is available at https://github.com/Rose-STL-Lab/CPTC.
@@ 1 Introduction
Uncertainty Quantification (UQ) is a key building block of reliable machine learning systems. Conformal prediction (CP) is a popular distribution-free UQ method that provides finite sample coverage guarantees without placing assumptions on the underlying data generating process (vovk2005algorithmic; lei2018distribution). Because of its simplicity and generality, conformal prediction has seen wide adoption in many tasks, ranging from healthcare angelopoulos2022conformal, robotics lindemann2023safe, to validating the factuality of generative models cherian2024large.
@@
State Space Models (SSMs) (kalman1960new; aoki2013state) are a powerful tool for modeling time series data, as they provide a structured framework to capture the underlying temporal dependencies through latent states. In particular, Switching Dynamical Systems (SDS) (ackerson1970state) is a type of SSM with an additional set of latent variables (known as switches) that represent the operating mode active at the current timestep. SDS introduces the inductive biases that allow SSMs to switch between a discrete set of dynamics, which can be learned from training data. It is shown to be a flexible, robust, and interpretable representation of time series with varying dynamics exhibiting change points (ansari2021deep; chen2022learning). SDS is deployed not only for energy karakatsani2008intra; song2014short; fan2008machine and traffic yu2004switching; cetin2006short forecasting as examples given above, but also on a wide range of applications including neuroscience glaser2020recurrent; rabinovich2006dynamical; kato2015global, engineering strachan2011measuring; wang2007actively, and sports analytics yamamoto2018switching, all would benefit from robust and calibrated uncertainty quantification.
@@ 2 Related Work
Probabilistic forecasting has become central to modern time-series analysis by modeling a distribution over future outcomes rather than a single-point forecast (benidis2022deep). Classic approaches often adopt Bayesian or approximate Bayesian strategies for uncertainty quantification: for instance, Bayesian Neural Networks (BNNs) (wang2019deep; wu2021quantifying; gal2017concrete) or neural process models (garnelo2018conditional; sun2024data) use neural networks to parameterize the underlying stochastic process. Other methods, such as DeepAR (salinas2020deepar), PatchTST (nie2023patchtst), or Temporal Fusion Transformer (lim2021temporal), generate probabilistic forecasts directly by minimizing a pinball loss (i.e., quantile loss), thereby learning predictive distributions.
A key challenge arises when change points occur - i.e., abrupt distribution shifts in the time series (montanez2015inertial). Change point detection and segmentation have been studied extensively (beyond our scope, see (aminikhanghahi2017survey) for a recent survey). Modeling time series with change points typically features state-space models (SSM), and in particular switching dynamical systems (SDS) (johnson2016composing; ansari2021deep; liu2023graph). These SDS approaches use variational inference to fit neural parameterizations of transitions and emission distributions. However, existing probabilistic forecasters often lack strict calibration guarantees and can be overconfident in practice (lakshminarayanan2017simple), and many require specialized architecture tuning for different tasks, limiting their applicability to real-world UQ challenges.
@@
Conformal prediction (CP) (vovk2005algorithmic) has emerged as a leading framework for UQ due to its algorithmic simplicity, broad applicability, and finite-sample coverage guarantees; see angelopoulos2021gentle for a comprehensive introduction. For time-series forecasting, some works focus on joint coverage over fixed-length horizons (Schaar2021conformaltime; sun2023copula; zhou2024heterogeneous); others assume stable underlying dynamics and adapt only to shifts in residual distributions (xu2021conformal) . tibshirani2019conformal uses known covariate shifts (by propensity weighting) to achieve theoretical coverage, but can fail in settings where the shift mechanism is unknown, which is typical for time series forecasting. hajihashemi2024multi introduces a multi-model ensemble CP framework, optimizing model selection within a pool of experts for combined coverage. CP has also been extended for change-point detection in time series via conformal martingales (vovk2021retrain; volkhonskiy2017inductive).
Recent works on online conformal prediction develop online strategies that guarantee marginal coverage for adversarial or nonstationary data streams. ACI (gibbs2021adaptive), AgACI (zaffran2022adaptive), DtACI (gibbs2024conformal), and Conformal PID control (angelopoulos2024pid) leverage online optimization and have a ”reaction” period when distribution shifts. Multivalid Prediction (MVP) (bastani2022practical) provides group coverage guarantees and learns a calibration threshold online, although it can require longer calibration windows. Closer to our work, SPCI (xu2023sequential) learns an autocorrelative estimation of the residual’s quantiles through Quantile Random Forest; whereas HopCPT (auer2023hopcpt) condition their conformal interval on similar parts of the time series using a Modern Hopfield Network. The drawback of these two methods is that the regression models are trained during inference time, which (1) requires a long cold start window to learn meaningful associations and (2) is computationally expensive. Our contribution is a CP algorithm that is computationally light during inference time, and can leverage state prediction capabilities within the SDS model to anticipate and adapt quickly to distribution changes.
@@ 3.1 Conformal Prediction
We say a sample is covered if the true value lies in the prediction interval $y^{n+1}\in\Gamma^{1-\alpha}(x^{n+1})$ Conformal prediction is guaranteed to produce valid prediction intervals (vovk2005algorithmic) if the calibration data and test data are exchangeable (in a dataset $\{(x^{i},y^{i})\}_{i=1}^{n}$ of size $n$, any of its nn! permutations are equally probable).
@@ 3.2 Switching Dynamical Systems for Modeling Time-series with Change Points
There have been many different variations and implementations of SDS. Classic SDS have memory-less state transitions, i.e. $p(z_{t}|z_{t-1})$ is parameterized by a stochastic transition matrix $\mathbf{A}\in\mathbb{R}^{K\times K}$. This results in an “open loop” system and the state duration follows a geometric distribution. Recent works have sought to fix this issue, for example by closing the loop using recurrent dependencies on $s_{t}$ (linderman2016recurrent), or modeling explicit durations (ansari2021deep). Our generative model is illustrated in figure 2): for generality, we factor out any non-Markovian covariates to model parameters $\theta$. The joint distribution can be factorized as
@@ 17 $\Gamma_{t}(x_{t})\leftarrow$ Aggregate the $\Gamma_{z,t}(x_{t})$s by Eqn 10 or Eqn 11;
The exchangeability assumption does not apply to the online setting of our work. If we naively calibrate with online data, when the data distribution changes, the coverage probability may deviate from the target level $1-\alpha$. Therefore, we follow online CP methods such as ACI (gibbs2021adaptive) to address this problem by continuously updating $\alpha_{z,t}$ to track a surrogate miscoverage rate $\alpha_{z,t}^{*}$ as defined in eqn 8 (for each state $z$ in our setting), an internal estimate of the target coverage.
@@
In eqn 8 $M_{z,t}$ measures the probability that the true label $y_{t}$ falls outside the prediction set $\Gamma_{z,t}$ when the predicted state is $z$. We track $\alpha^{*}_{z,t}$ over time by updating the estimate $\alpha_{z,t}$ with online optimization. In our implementation, we use the simple update from ACI (Eqn 9), though more sophisticated approaches such as (angelopoulos2024pid) may be used in its place without affecting the overall algorithm.
@@
Our algorithm can be viewed as running multiple instances of online conformal inference (lines 14-16 of algorithm 1), one for each underlying state, and adaptively aggregating the confidence intervals when the underlying mode is switching. This differs from the setting where the temporal correlation between residuals needs to be learned online (gibbs2024conformal; bastani2022practical; xu2023sequential; auer2023hopcpt); the state model allows us to leverage training data for such information. Our algorithm is modular - the state model, forecaster model, and the online adaptive conformal prediction algorithms all operate independently of each other. For example, the state-specific online coverage tracking (line 15 of algorithm 1) can be replaced with other online conformal prediction variants for data-specific applications. Our theoretical guarantees are agnostic to the forecaster models used.
@@ 5 Experiments
Table 1: Performance on synthetic and real-world datasets with target confidence $1-\alpha=0.9$ (for horizon $T=200$ for synthetic datasets, 300300 for electricity and traffic, and $60$ for bee, mean $\pm$ standard deviation of the test samples). Methods that are invalid (coverage below $90\%$) are grayed out. Our method achieves a high level of calibration (coverage is close to $90\%$) consistently.
@@
We selected the following baseline methods. RED-SDS (Recurrent Explicit Duration Switching Dynamical System) ansari2021deep is a state-of-the-art Bayesian model that learns state transitions (implementation details in Appendix B.1). We also use REDSDS as the base predictor for mode switching for our method in the experiments. CP is a generalization of conformal prediction to the online setting, also known as Online Sequential Split Conformal Prediction in zaffran2022adaptive. We choose ACI (gibbs2021adaptive) to represent online CP algorithms leveraging online optimization. SPCI (xu2023sequential) and HopCPT auer2023hopcpt are state-of-the-art CP algorithms that learns adaptive predictive intervals by quantile regression and Modern Hopfield Networks learned from the time series respectively.
@@
For real-world datasets, the Electricity and Traffic datasets from (Dua2017UCI) have hourly frequency and exhibit seasonality both in terms of the time series itself and volatility. The honey bee trajectory dataset (oh2008learning) is the most complex, composed of 4-dimensional trajectories with length averaging to 900 frames, where the bees’ dance can be decomposed into “left turn”, “right turn” and “waggle”.
@@ 5.1 Analysis of Coverage and Sharpness Results
Results for uncertainty quantification in table 4 show that we achieve significantly better validity on all datasets compared to baselines. RED-SDS’s intervals fail due to lack of calibration, a common pitfall of probabilistic methods (shown also in xu2023sequential). CP’s under-coverage shows that the data are not exchangeable. In the presence of change points, ACI’s coverage fluctuates dramatically and did not converging within the horizon (more visualizations in Appendix B.5). For SPCI and HopCPT, which also guarantees asymptotic validity, the invalidity is likely because they are designed for large datasets ($T\geq 10,000$ for HopCPT), and did not have enough data to react or learn useful residual patterns for the relatively short prediction horizon of our datasets ($T=200$ for synthetic datasets, 300 for electricity and traffic, and 60 for bees). Figure 3 provides a qualitative example of how the prediction intervals compare across different methods. Our method (purple shaded regions) does not over-cover during nonvolatile hours, nor under-cover during busy ones, as does ACI.
@@ A.2 Proofs: Lemma
We derive the following lemma from the convergence result of Proposition 4.1 in [gibbs2021adaptive] with two additional assumptions also according to [gibbs2021adaptive]: (1) the existence of a single fixed optimal target $a^{*}\in[0,1]$, and (2) $\mathbb{E}[err_{t}|a_{t}]=M(a_{t})$. These are reasonable assumptions that allow us to quantify distribution shifts and are necessary for analyzing adaptation behavior (whereas proposition 4.1 itself is a worst-case upper bound.) We refer the readers to the original paper for the complete derivation.
@@ B.3 Dataset descriptions, continued
Electricity and Traffic. The Electricity and Traffic dataset are from the UCI machine learning repository [Dua2017UCI] and we use the train/test split from GluonTS [alexandrov2020gluonts]. The electricity dataset consists of hourly electricity consumption data for 370 customers, each with 5,833 observations. The traffic dataset consists of hourly occupancy rates from 963 road sensors, each with 4,000 data points.
Dancing Bees. The honey bee trajectory dataset is from [oh2008learning]. Honey bees convey information about the location and distance to a food source through a dance carried out inside the hive. This dance consists of three distinct phases: “left turn”, “right turn”, and “waggle”. The duration and orientation during the waggle phase represent the distance and direction to the food source. Experiments are conducted on 6 trajectories, with lengths of 1058, 1125, 1054, 757, 609, and 814 frames, respectively. In this paper, the prediction interval is on the first two dimensions (the coordinates of the bee) only, but all dimensions are used for prediction.
@@ B.4 Additional Experiments
We add comparison to three more baselines that handles distribution shift in time series data. They are DtACI [angelopoulos2024online], ECI [wu2025error], and L-ARC [zecchin2024localized].
@@
Table 4: Performance on synthetic and real-world datasets with target confidence $1-\alpha=0.9$ (for horizon $T=200$ for synthetic datasets, 300300 for electricity and traffic, and $60$ for bee, mean $\pm$ standard deviation of the test samples). Methods that are invalid (coverage below $90\%$) are grayed out. Our method achieves a high level of calibration (coverage is close to $90\%$) consistently.
@@
CPTC offers distinct advantages over SPCI and HopCPT in data-limited regimes and applications requiring rapid deployment. While SPCI and HopCPT achieve excellent long-term performance on extensive time series (T $\geq$ 10,000) by learning residual patterns through quantile regression and Hopfield networks respectively, they require substantial data to train these models effectively during inference. Our results demonstrate that CPTC significantly outperforms these methods on shorter horizons (T $\leq$ 300), where insufficient data prevents SPCI and HopCPT from learning meaningful temporal correlations or residual patterns.
@@
We designed the number of latent states to be a hyperparameter given by the user. Our method ensures that the validity guarantee still holds even if the state cardinality is incorrect (theorem 4.2), as incorrect state cardinality can be formulated as a case of incorrect classification. Determining state space cardinality is a well-studied problem (see for example [branicky1994stability]), and we consider it to be beyond the scope of our paper.